How can I simplify this with all positive exponents?:
5x^-1 / y^-1

Found 2 solutions by Fombitz, polas4sure@gmail.com:
Answer by Fombitz(32388)
Answer by polas4sure@gmail.com(2)

Simplify: 5x^-1 / y^-1
SOLUTION.
5x^-1/y^-1
= 5 (x)^-1/1/y

= (5/x) / (1/y)
= (5/x) * (y/1)
therefore